A sports training camp has 60% boys. 20 random players win a trophy each day. 68% of the winners can be expected to be between ---- boys

A> 59.8905% and 60.1095%
B> 49.05% and 70.95%
C> 59.8658% and 60.1342%
D> 46.58% and 73.42%
E> None of the above
1
Expert's answer
2020-03-20T17:57:09-0400

We need to construct the 68% confidence interval for the population proportion. We have been provided with the following information about the sample proportion: "\hat{p}=0.60, n=20."

The critical value for "\alpha=0.32" is "z_c=z_{1-\alpha/2}=0.994." The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{{\hat{p}(1-\hat{p}) \over n}},\hat{p}+z_c\sqrt{{\hat{p}(1-\hat{p}) \over n}})"

"CI(Proportion)=(0.6-0.994\sqrt{{0.6(1-0.6) \over 20}},0.6-0.994\sqrt{{0.6(1-0.6) \over 20}})\approx"

"\approx(0.491, 0.709)"

B) 49.05% and 70.95%